OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (5, 6) and (9, 14)? y = x + @bruhhh

OpenStudy (anonymous):

Okay so this has the same idea as the last one, except the slope isn't given. So you've gotta calculate it by using this: \[\frac{ y2-y1 }{ x2-x1 }\] Plugging those in will make it\[\frac{ 14-6 }{ 9-5 }\] So evaluate your slope first, tell me what you get :)

OpenStudy (anonymous):

I'm also doing this do you mind if I stay and learn too cause I've had similar questions and im not good at them

OpenStudy (anonymous):

8/4 @bruhhh and for you @shinebrightlikeadimon no i don't mind you can stay

OpenStudy (chosenmatt):

hey @TeenWolfGirl im gonna help you one sec

OpenStudy (anonymous):

@shinebrightlikeadimon yeah i dont mind either :) haha So your slope can simplify even more, whats 8/4?

OpenStudy (anonymous):

ok and @bruhhh why did i get a medal?

OpenStudy (anonymous):

4/2 and if you simplify that you get 2/1 which = 2

OpenStudy (anonymous):

i gave you one ^.^ so since the slope is 8/4 m = 2?

OpenStudy (anonymous):

Lol whaaa, I don't even know how to give medals so no idea

OpenStudy (anonymous):

oh!!! @shinebrightlikeadimon gave me one i still don't know why?

OpenStudy (anonymous):

why in the world am i getting a medals i still don't know the answer

OpenStudy (anonymous):

Yesss @shinebrightlikeadimon So now you guys need to use point slope form using one of the ordered pairs, which is: y-y1=m(x-x1) Plugging in either ordered pair will get you the same answer but I'm gonna use (5,6) since the numbers are smaller lol So it will look like: y-6=2(x-5) Do ya'll know what to do from there to get it in y=mx+b form?

OpenStudy (anonymous):

i only got 2

OpenStudy (anonymous):

*looks at @bruhhh*

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait do you simplify ?

OpenStudy (anonymous):

yea that's how you get the answer

OpenStudy (anonymous):

so simplify and add 6 to both sides?

OpenStudy (anonymous):

*smiles* sweetheart i'm not that smart no mean to be rude but if i was i wouldn't need this website to teach me

OpenStudy (anonymous):

Yes! @shinebrightlikeadimon @TeenWolfGirl do you get that? It's the same thing we did in the last problem, I'll type the example in a sec

OpenStudy (anonymous):

thnx ? @TeenWolfGirl and great cause i wasn't sure @bruhhh

OpenStudy (anonymous):

So we've got y-6=2(x-5) You've gotta distribute the slope (2) and that's what we mean by simplify, so it'll look like y-6=2x-10 Then add the 6 to both sides and what's your answer @TeenWolfGirl :)

OpenStudy (anonymous):

@shinebrightlikeadimon no prob! :)

OpenStudy (anonymous):

y=8x-16

OpenStudy (anonymous):

y - 6 = 2x - 10 + 6 + 6 y = 2x - 4 ?

OpenStudy (anonymous):

the slope is 2 not 8 @TeenWolfGirl

OpenStudy (anonymous):

how i though you add 10+6 not subtract?

OpenStudy (anonymous):

Sorry, I've gotta get in the shower and go after this one but I'll be on later if you need more help. But yeah. @shinebrightlikeadimon is right. You are adding 6, but the equation says the 10 is negative. So it's -10+6 :)

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

so its y=2x+2?

OpenStudy (anonymous):

Here use a calculator for that type of addition if you need it, it'll help a ton! -10+6 is -4 So it's y=2x-4 Make sense?

OpenStudy (anonymous):

hehe yea forgot about that thx

OpenStudy (anonymous):

geez i want to give both medal but i can only give one @bruhhh can you give @shinebrightlikeadimon a medal?

OpenStudy (anonymous):

Lol okiedoke, I'll be on later ^.^ Good luck!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

to both of you

OpenStudy (anonymous):

np and thanks you both

OpenStudy (anonymous):

hehe np